OpenStudy (anonymous):

The cafe charges $3.50 for a buffet lunch of 2 eggs and 3 strips of bacon. The charge is $ 5.00 for 3 eggs and 4 strips of bacon . Find the cost of each egg and the cost of each strip of bacon?

OpenStudy (jack1):

hiya @cutegirl this looks like simultaneous equations to me, mind if i have a go?

OpenStudy (anonymous):

go ahead

OpenStudy (jack1):

350 = 2e + 3b EQN 1 500 = 3e + 4b EQN 2 => \[\large EQN~1 \times 3 = 3(350) = 3(2e + 3b) = EQN~1a\] \[\large 1050 = 6e + 9b = EQN~1a\] \[\large EQN~2 \times 2 = 2(500) = 2(3e + 4b) = EQN~2a\] \[\large 1000 = 6e + 8b = EQN~2a\]

OpenStudy (jack1):

so as u can see, we now have 6e in both equations so rearrange the equations into the form: "6e = ...?" then you can solve

OpenStudy (jack1):

do u think u'll be ok from here, or would u like me to keep going?